1) Convert the following polar equation of r=2sintheta into rectangular form.

2) List and find all of the sixth roots of 4.
Answer #1

. 80 = 2 sino r²= =rcoso = 2rsino Y=rsino (x²+ y²) = ay + 14-45% = 1 - (CIRCLE) 2. het to=4 26= 4 (cos Qka tisin 2kt) i(2x) x